# A Comprehensive Evaluation on Quantization Techniques for Large Language Models

### v1

@@ Abstract
For large language models (LLMs), post-training quantization (PTQ) can significantly reduce memory footprint and computational overhead. Model quantization is a rapidly evolving research field. Though many papers have reported breakthrough performance, they may not conduct experiments on the same ground since one quantization method solution usually contains multiple components. In addition, it is also important to analyze the theoretical connections among numerous existing methods to gain in-depth understanding in this field. To bridge these gaps, in this paper, we timely conduct an extensive review of the state-of-the-art methods and also perform comprehensive evaluations on the same ground to ensure fair comparisons of these methods. To our knowledge, this fair and extensive investigation remains critically important yet significantly underexplored in the research community.
To better understand the theoretical connections, we decouple the published quantization methods into two steps: pre-quantization transformation and quantization error mitigation. We define the former as a preprocessing step applied before quantization to reduce the impact of outliers, making the data distribution flatter and more suitable for quantization. Quantization error mitigation, on the other hand, involves techniques that offset the errors introduced during quantization, thereby enhancing model performance. Meanwhile, we extensively evaluate and analyze the impact of different components of quantization methods. Additionally, we also timely analyze and evaluate the latest MXFP4 data format and its performance. Our experimental results demonstrate that the optimized rotation and scaling yield the best performance for pre-quantization transformation, and combining low-rank compensation with GPTQ occasionally outperforms using GPTQ alone for quantization error mitigation. Furthermore, we explore the potential of the latest MXFP4 quantization and reveal that the optimal pre-quantization transformation strategy for INT4 does not generalize well to MXFP4, inspiring further investigation in this field.
@@ 1 Introduction
Post-training quantization (PTQ) is a promising solution to reduce the memory and computational cost of LLMs, enhancing their inference efficiency. LLMs quantization has been studied in a wide spectrum of precision configurations. Representative settings include W4A4 [1, 2, 3, 4], W8A8 [5, 6, 7] and W4A16 [8, 9, 10, 11]: W4A4 quantizes both weights and activations to 4-bit precision (INT4); W8A8 and W4A16 quantize weights to INT8 and INT4 and activations to INT8 and FP16/BF16 respectively. Among these 3 settings, W4A4 is the current research focus because it provides the best performance in terms of both memory efficiency and inference speed, with the greatest research significance.
However, despite the wide range of methods explored in previous studies, W4A4 quantization still struggles to achieve lossless performance, especially for smaller LLMs, such as those with 1B, 3B, and 7B parameters, which often results in increased perplexity and degraded performance on various tasks. Furthermore, LLaMA3 experiences a larger performance drop after quantization compared to LLaMA2 when using many popular quantization methods [3, 4, 12], facing a greater demand for better W4A4 quantization methods. Therefore, W4A4 quantization for LLMs still remains a challenging task.
To review and evaluate the state-of-the-art quantization methods, in particular, under the setting of W4A4, we find two main characteristics of these methods. First, existing methods are often combinations of multiple techniques, since using a single technique alone typically fails to achieve excellent quantization performance. In SpinQuant [3], after applying trained rotations to adjust the model, the GPTQ [8], a gradient-based quantization error mitigation method is further applied to quantize the rotated weights. Similarly, OSTQuant [8] employs multiple techniques to adjust the weights and activations, including rotation and scaling. At the same time, there are many weight-only quantization methods, some of which can be combined with pre-quantization transformation techniques and extended to weight-activation quantization. AWQ [11] shares similar pre-quantization transformation (scaling) with SmoothQuant [6], even though their motivations differ. Also, QuIP# [10] shares similar pre-quantization transformation (random hadamard rotation) with QuaRot [13]. Apart from the methods that involve both weight and activation transformations, there are also approaches that are orthogonal to pre-quantization transformation that can be used to compensate for the loss between quantized models and full precision models, such as GPTQ [8] and low-rank branches computed from quantization error [14, 15, 16]. Second, quantization involves a variety of configurations, such as granularity and symmetric or not. Different settings can lead to different performance outcomes and often involve trade-offs between efficiency and accuracy. For different precision levels, the optimal quantization settings may vary. Evaluating the performance under various quantization configurations for W4A4 quantization is of great importance.
@@
In this paper, as aforementioned in the first characteristic, the existing quantization solutions are very complicated and they usually consists of many algorithms. To make our evaluation on the same ground and achieve a better understanding of existing methods, we decompose the quantization process into two key steps: pre-quantization transformation and quantization error mitigation. The pre-quantization transformation step refers to preprocessing the data before quantization in order to mitigate the effect of outliers making the data distribution smoother and more quantization-friendly. The quantization error mitigation step involves applying techniques to compensate for the errors introduced by quantization, thereby reducing the overall quantization error and improving performance. For each step, we select several state-of-the-art techniques and conduct an evaluation of the effectiveness of their combinations. Second, existing solutions usually do not use exactly the same setting, e.g. granularity, leading to unfair comparison. To bridge this gap, we conduct a comprehensive evaluation on the same quantization configurations and provide recommendations on selecting appropriate configurations. And we integrate the methods used in this paper in our code library to facilitate performance evaluation.
@@
Furthermore, the NVIDIA’s GeForce RTX 50 Series latestly released based on the new RTX Blackwell architecture includes support for the FP4 data format, including NVFP4 and MXFP4, in its new Tensor Cores, which show great advantages over INT4. However, the research on FP4 has not been well explored. To bridge this gap, we conduct research on weight-activation FP4 quantization using the popular methods in INT4, showing the great timeliness of our research.
Methods Data Normalization Error Compensation
SmoothQuant [6] scale -
Outlier Suppression+ [7] scaling + shifting -
@@
OmniQuant [2] optimized scaling + shifting GPTQ
@@
SpinQuant [3] optimized rotation GPTQ
@@
FlatQuant [12] optimized scaling + rotation GPTQ
OSTQuant [4] optimized scaling + rotation GPTQ
@@
2 Review on Existing Quantization Methods
In this section, we first review recent quantization algorithms in Section 2.1 and Section 2.2 in terms of two steps: pre-quantization transformation and quantization error mitigation. Table 1 shows our decomposition of the popular quantization methods. Section 2.3 and Section 2.4 focus on settings, including symmetricity and granularity of quantization methods. Meanwhile, Table 2 summarizes the quantization settings reported in the experimental results of these methods.
@@ 2 Review on Existing Quantization Methods
Table 2: Symmetry and granularity of quantization methods. In granularity, T stands for per-tensor, C represents per-channel or per-token, and G means per-group
2.1 Pre-quantization transformation
Figure 1: Pre-quantization transformation: Transforming activations and weights to render them more easier to quant
2.1.1 Shifting
Activations distribution is often asymmetric, and this asymmetry varies across different channels. Aligning the center of each channel can improve the performance of quantization, since we quantize activations per-token rather than along the channel dimension. Outlier Suppression+ [7] proposes shifting activation values to eliminate the asymmetry feature across channels. The shifting process is to use a vector that transfer the channels of the activation. For a linear layer, it is formulated as:
$Y=XW+B=(\underbrace{(X-T)}_{\hat{X}})W+(\underbrace{TW+B}_{\hat{B}})$ (1)
where the layer input $X\in\mathbb{R}^{T\times C_{in}}$ is shifting and becomes $\hat{X}$. $T\in\mathbb{R}^{1\times C_{in}}$ is a vector representing the channel-wise shifting parameters. To keep mathematical equivalent, an additional term, the product of $T$ and $W\in\mathbb{R}^{C_{in}\times C_{out}}$, is added to the bias $B\in\mathbb{R}^{1\times C_{out}}$. Outlier Suppression+ [7] uses calibration to get the vector $T$, which is computed channel-wise as the average of its minimum and maximum values:
$t_{i}=\frac{\min(X_{i})+\max(X_{i})}{2}$ (2)
where $t_{j}$ denotes the $j$-th element of $T$, corresponding to the $j$-th channel. $X_{j}$ represents the $j$-th input channel of $X$. OmniQuant [2] employs learnable channel-wise shifting for outlier suppression, combined with the learnable scaling technique discussed in Section 2.1.2, and jointly optimize these two kinds of pre-quantization transformation parameters.
2.1.2 Scaling
Scaling is a commonly used pre-quantization transformation method, which scales activations in different input channels before quantization, and inversely scales the corresponding weight channels to make the results stay unchanged. SmoothQuant [6] first propose using scaling to smooth the quantization difficulty from activations to weights. Compared to weights, activations have more outliers that make quantization harder. It is observed that outliers in activations appear in certain channels, so applying channel-wise scaling to activations can eliminate the outliers to a certain degree, using the following equation for linear layers:
@@ 2.1.2 Scaling
where the weights of a linear layer $W$ and the layer input $X$ is scaling and becomes $\hat{X}$ and $\hat{W}$. $S$ is a diagonal matrix composed by scaling factors of different channels, which can be obtained in various ways. SmoothQuant [6] uses calibration to compute $S$. The calibration process can be formulated as:
$s_{j}=max(|X_{j}|)^{\alpha}/max(|W_{j}|)^{1-\alpha}$ (4)
where $X_{j}$ and $W_{j}$ denote the $j$-th input channel of $X$ and $W$, respectively. $s_{j}$ is the $j$-th diagonal element of $S$, corresponding to the $j$-th channel. $\alpha$ is a hyperparameter that controls the degree to which outliers are transferred between activations and weights.
AWQ [11], as a weight-only quantization method, approaches the problem from the perspective of protecting salient weights, applying activation-aware scaling. Unlike adopting the hand-crafting parameters in calibration, the scaling operation in OmniQuant [2] uses learnable scaling, where $S$ is learnable parameters that are optimized with block-wise quantization error minimization.
FlatQuant [12] and OSTQuant [4] both use learnable per-channel scaling, but at different stages of pre-quantization transformation, leading to different practical meanings. Both perform a diagonal scaling operation, yet FlatQuant scales before rotation, allowing its scaling factors to be initialized via Eq. (4), functionally similar to SmoothQuant. In contrast, OSTQuant applies scaling after rotation, meaning this diagonal scaling is a post-rotation adjustment rather than smoothing the original channels.
2.1.3 Rotation
The rotation operation is a technique that adjusts data distribution by multiplying the data matrix with an orthogonal matrix. This operation largely changes the activation distributions, reduces outliers in the data and thereby enhancing quantization accuracy. The rotation process in a linear layer can be formulated as:
@@ 2.1.3 Rotation
QuIP [9] first introduced rotation with random orthogonal matrices to reduce incoherence in weight-only quantization. QuIP# [10] replaces this with random Hadamard matrices for greater efficiency. QuaRot [13] extends rotation to both weights and activations quantization using random Hadamard matrices, effectively eliminating outliers. SpinQuant [3] further improves stability by optimizing the rotation matrix on the Stiefel manifold to minimize quantization-induced performance variance, while still defaulting to random Hadamard for a few online rotations.
@@
Rotation is a highly effective pre-quantization transformation technique for reducing outliers in low-bit quantization, and it has also been widely adopted in many other quantization methods (e.g., ResQ [19], QServe [17], FlatQuant [12] and OSTQuant [4], etc), which combine rotation with other pre-quantization transformation techniques to further improve performance.
2.2 Quantization error mitigation
@@
The quantization error mitigation step employs techniques that mitigate the error introduced by quantization by compensating them. This error between the original model and the quantized model can be computed from calibration datasets, and be compensated to the original weights in linear layers or through low-rank branches. Fig. 2 demonstrates two commonly used quantization error mitigation methods.
Figure 2: Quantization error mitigation: compensate the error produced by quantization. (1) Self-compensation [8, 9] calculates weight compensation terms through quantization-induced loss errors; (2) Low-rank branch compensation [14, 16, 18, 15] constructs auxiliary compensation pathways via matrix factorization of post-quantization weight errors.
2.2.1 Self compensation
GPTQ [8] is a quantization method for weight quantization that uses Hessian-based quantization error mitigation. GPTQ builds on the OBQ [20] method that solves the error introduced by quantization and the optimal weight perturbation. For weight quantization, the loss error introduced by quantization can be defined as:
$\epsilon=\mathcal{L}(W_{q})-\mathcal{L}(W)=\mathcal{L}(W+\delta)-\mathcal{W}$ (6)
where $L(W)$ is the loss function of the model evaluated at weights $W$, and $\delta$ represents the weight perturbation. Using a Taylor expansion, we can get the Hessian-based approximation of the quantization loss, which can be formulated as:
$\epsilon=\nabla\mathcal{L}(W)^{\top}\delta+\frac{1}{2}\delta^{\top}H(W)\delta+\mathcal{O}(\left\|\delta\right\|^{3})\approx\frac{1}{2}\delta^{\top}H(W)\delta$ (7)
Unlike OBQ that uses greedy order to quantize weights, GPTQ quantized all rows of weights in a fixed order. The GPTQ algorithm performs quantization column by column in blocks of size B. For each column within the block, the quantization error is computed when quantized, and then compensates on the remaining weights. Once all columns of a block is quantized, the remaining weights in the matrix are also updated.
QuIP [9] also minimizes a quadratic proxy of the weight error to implement an optimal adaptive rounding procedure called LDLQ. In this method, an LDL decomposition of the Hessian matrix is performed to construct a feedback matrix $U$, which is then used to form a compensation term that corrects for the quantization-induced error. In QuIP, it is proven that for Eq. (7), performing an LDL decomposition on the Hessian matrix yields an optimal choice of $U$, and GPTQ is a special case of LDLQ.
2.2.2 Low-rank branch compensation
@@
Beyond compensating errors directly on the original weights, an alternative approach is to introduce an auxiliary branch dedicated to error correction. Some prior works leverage a low-rank term to reconstruct quantization error (e.g., ZeroQuant-v2 [14], CALDERA [16], LQER [18], QERA [15], etc). After incorporating a low-rank branch, the output $Y$ is computed as:
$Y=XW=XW_{q}+X\Delta W\approx XW_{q}+XAB$ (8)
@@
where $\Delta W=W-W_{q}$ denotes the weight reconstruction error introduced by quantization. $A\in\mathbb{R}^{C_{in}\times r}$, $B\in\mathbb{R}^{r\times C_{out}}$ are learnable low-rank matrices, and $r\ll\min(d,k)$ is the bottleneck rank. The product $AB\in\mathbb{R}^{d\times k}$ serves as a low-rank approximation of the quantization error.
ZeroQuant-v2 [14] and LQER [18] reconstruct the weight quantization error through SVD-based low rank approximation. LQER computes the low rank matrices $A,B$ by using SVD composition to the quantization error matrix $\Delta W$:
$\Delta W=U\Sigma V^{T}\approx\underbrace{U_{k}}_{A_{k}}\cdot\underbrace{\Sigma_{k}V_{k}^{T}}_{B_{k}}=A_{k}B_{k}$ (9)
where $U\in\mathbb{R}^{m\times n}$, $V\in\mathbb{R}^{n\times n}$, $\Sigma\in\mathbb{R}^{m\times n}$ and $U_{k}\in\mathbb{R}^{m\times k}$, $V_{k}\in\mathbb{R}^{n\times k}$, $\Sigma_{k}\in\mathbb{R}^{k\times k}$.
@@
Based on LQER, L2QER, which is proposed in LQER, scales the $\Delta W$ before applying SVD, and inversely apply scales in $A_{k}$. This operation aims to more precisely approximated the quantization error corresponding to the salient weights. The process can be formulated as:
$\Delta\hat{W}=S\Delta W=U\Sigma V^{T}\approx\underbrace{U_{k}}_{A_{k}}\cdot\underbrace{\Sigma_{k}V_{k}^{T}}_{B_{k}}=A_{k}B_{k}$ (10)
$\hat{A_{k}}=S^{-1}A_{k},\hat{B_{k}}=Bk$ (11)
@@
In addition to the aforementioned approaches, CALDERA [16] and QERA [15] provide analyses from different perspectives, deriving the optimal low-rank terms for quantization error reconstruction. In contrast to approaches that introduce low-rank branches externally, QwT [21] integrates lightweight linear layers within each decoder layer block of LLMs, directly compensate quantization errors, offering new insights for quantization error mitigation.
2.3 Symmetric vs. asymmetric

### v2

@@ Abstract
For large language models (LLMs), post-training quantization (PTQ) can significantly reduce memory footprint and computational overhead. Model quantization is a rapidly evolving research field. Though many papers have reported breakthrough performance, they may not conduct experiments on the same ground since one quantization method solution usually contains multiple components. In addition, it is also important to analyze the connections among numerous existing methods to gain in-depth understanding of this field. To bridge these gaps, in this paper, we timely conduct an extensive review of the state-of-the-art methods and also perform comprehensive evaluations on the same ground to ensure fair comparisons of these methods. To our knowledge, this fair and extensive investigation remains critically important yet significantly underexplored in the research community.
To better understand the connections, first, we decouple the published quantization methods into two steps: pre-quantization transformation and quantization error mitigation. We define the former as a preprocessing step applied before quantization to reduce the impact of outliers, making the data distribution flatter and more suitable for quantization. Quantization error mitigation, on the other hand, involves techniques that offset the errors introduced during quantization, thereby enhancing model performance. Second, we extensively evaluate and analyze the impact of different settings of quantization methods, including granularity and symmetry. Third, we also timely analyze and evaluate the latest MXFP4 and NVFP4 data format and its performance.
Our experimental results first demonstrate that the optimized rotation and scaling yield the best performance for pre-quantization transformation, and combining low-rank compensation with GPTQ occasionally outperforms using GPTQ alone for quantization error mitigation. Second, as the granularity becomes finer, the performance improves, but this also incurs additional storage overhead. We also verified that applying asymmetric quantization to weights yields much smaller benefits compared to applying it to activations, which is consistent with the common practice of using asymmetric quantization primarily to activations. Third, we explore the potential of the latest FP4 quantization and find that the format and precision of scaling factors have a significant impact on the performance. We also show the performance of the optimal pre-quantization transformation strategy for INT4 on FP4, revealing that rotation-based methods demonstrate little improvement in performance for MXFP4 and NVFP4, inspiring further investigation in this field.
@@ 1 Introduction
Post-training quantization (PTQ) is a promising solution to reduce the memory and computational cost of LLMs, enhancing their inference efficiency. LLMs quantization has been studied in a wide spectrum of precision configurations. Representative settings include W4A4 [1, 2, 3, 4], W8A8 [5, 6, 7] and W4A16 [8, 9, 10, 11]: W4A4 quantizes both weights and activations to 4-bit precision (INT4); W8A8 and W4A16 quantize weights to INT8 and INT4 and activations to INT8 and FP16/BF16 respectively. Among these three settings, W4A4 is currently the main focus of research, as it holds great potential for improving memory efficiency and inference speed. However, applying lossless W4A4 quantization to LLMs remains a challenging problem. Thus, we concentrate our efforts on exploring W4A4 quantization.
@@
In this paper, we mainly investigate three dimensions. First, we find that existing methods are often combinations of multiple techniques, since using a single technique alone typically fails to achieve excellent quantization performance. OSTQuant [8] and FlatQuant [12] employ multiple techniques to adjust the weights and activations, including rotation and scaling, and also apply the GPTQ [8], a gradient-based method for mitigating quantization error, to quantize the transformed weights. Besides, many techniques share similar transformations, for instance, Quip# [10] and QuaRot [13] both use rotation, also, AWQ [11] and SmoothQuant [6] both use scaling. To achieve a better understanding of existing methods and make our evaluation on the same ground, we decompose the quantization process into two key steps: pre-quantization transformation and quantization error mitigation. The pre-quantization transformation step refers to preprocessing the data before quantization in order to mitigate the effect of outliers, making the data distribution smoother and more quantization-friendly. The quantization error mitigation step involves applying techniques to compensate for the errors introduced by quantization, thereby reducing the overall quantization error and improving performance. For each step, we select several state-of-the-art techniques and conduct an evaluation of the effectiveness of their combinations.
Second, quantization involves a variety of configurations, such as granularity and whether symmetric or not. Different settings can lead to different performance outcomes and often involve trade-offs between efficiency and accuracy. For different precision levels, the optimal quantization settings may vary. And existing solutions usually do not use exactly the same setting, leading to unfair comparison. To bridge this gap, we conduct a comprehensive evaluation on the same quantization configurations and provide recommendations on selecting appropriate configurations. And we integrate the methods used in this paper into our code library to facilitate performance evaluation.
@@
Third, NVIDIA’s GeForce RTX 50 Series, which is newly released and based on the Blackwell architecture, includes support for the FP4 data format (e.g., NVFP4 and MXFP4) in its Tensor Cores, showing significant advantages over INT4. However, the research on FP4 has not been well explored. To bridge this gap, we conduct research on weight-activation FP4 quantization using the popular methods in INT4, showing the great timeliness of our research. For FP4 quantization, we also observe that the format and precision of scaling factors have a significant impact on the performance.
2 Review on Existing Quantization Methods
In this section, we first review recent quantization algorithms in Section 2.1 in terms of two steps: pre-quantization transformation and quantization error mitigation. Table 1 shows our decomposition of the popular quantization methods. Section 2.2 focuses on settings, including symmetry and granularity of quantization methods. Meanwhile, Table 2 summarizes the settings of existing methods, showing that they are not compared on the same ground. Section 2.3 discusses FP4 quantization and its variants, including MXFP4 and NVFP4.
Methods Pre-quantization Transformation Quantization Error Mitigation
SmoothQuant [6] Scaling RTN
Outlier Suppression+ [7] Shifting + Scaling RTN
@@ 2 Review on Existing Quantization Methods
OmniQuant [2] Shifting+Scaling GPTQ
@@
SpinQuant [3] Rotation GPTQ
@@
FlatQuant [12] Scaling + Rotation GPTQ
OSTQuant [4] Scaling + Rotation GPTQ
@@
2.1 Two-steps of quantization
2.1.1 Pre-quantization transformation
@@
Table 2: Symmetry and granularity of quantization methods. In granularity, T stands for per-tensor, C represents per-channel or per-token, and G means per-group.
Figure 1: Pre-quantization transformation: a process that transforms activations and weights to make them easier to quantize. The left graph illustrates a tensor that is difficult to quantize due to the presence of numerous outliers, while the right graph shows that after the pre-quantization transformation, the tensor becomes more uniformly distributed, making quantization easier.
Shifting
Activation distribution is often asymmetric, and this asymmetry varies across different channels. Aligning the center of each channel can improve the performance of quantization, since we quantize activations per-token rather than along the channel dimension. Outlier Suppression+ [7] proposes shifting activation values to eliminate the asymmetry feature across channels. The shifting process is to use a vector $T$ that transfers the channels of the activation. For a linear layer, it is formulated as:
[…]
where the layer input $X\in\mathbb{R}^{T\times C_{in}}$ is shifting and becomes $\hat{X}$. $T\in\mathbb{R}^{1\times C_{in}}$ is a vector representing the channel-wise shifting parameters. To keep mathematical equivalence, an additional term, the product of $T$ and $W\in\mathbb{R}^{C_{in}\times C_{out}}$, is added to the bias $B\in\mathbb{R}^{1\times C_{out}}$.
Methods Shifting Strategy Shifting parameters $T$
OS+ [7] Pre-scaling shifting $T_{i}=\frac{\min(X_{i})+\max(X_{i})}{2}$
OmniQuant [2] Pre-scaling shifting Gradient-based optimization
Table 3: Comparison of shifting-based quantization methods.
As shown in Table 3, Outlier Suppression+ [7] uses calibration to get the vector $T$, while OmniQuant [2] employs learnable channel-wise shifting for outlier suppression, combined with the learnable scaling technique discussed in the next paragraph, and jointly optimizes these two kinds of pre-quantization transformation parameters.
Scaling
Scaling is a commonly used pre-quantization transformation method, which scales activations in different input channels before quantization, and inversely scales the corresponding weight channels to make the results stay unchanged. SmoothQuant [6] first proposes using scaling to smooth the quantization difficulty from activations to weights. Compared to weights, activations have more outliers that make quantization harder. It is observed that outliers in activations appear in certain channels, so applying channel-wise scaling to activations can eliminate the outliers to a certain degree, using the following equation for linear layers:
@@
where $S$ is a diagonal matrix composed of scaling factors of different channels, which can be obtained in various ways.
Methods Scaling Strategy Scaling parameters $S$
SmoothQuant [6] Scaling only $\begin{aligned} &S_{i}=\frac{\max(|X_{i}|)^{\alpha}}{\max(|W_{i}|)^{1-\alpha}}\end{aligned}$
AWQ [11] Scaling only $\begin{aligned} &S_{i}=|X_{i}|_{avg}^{\alpha}\\ &\alpha^{*}=\arg\min_{\alpha}\mathcal{L}(S_{Xi}^{\alpha})\end{aligned}$
OS+ [7] Post-shifting scaling $\begin{aligned} &S_{i}=\max\left(1.0,\frac{\max(X_{i}-T_{i})}{t}\right)\\ &t^{*}=\arg\min_{t}\mathcal{L}(S_{i})\end{aligned}$
OmniQuant [2] Post-shifting scaling Gradient-based optimization
FlatQuant [12] Pre-rotation scaling Calibration initialization and gradient-based optimization
OSTQuant [4] Post-rotation scaling Gradient-based optimization
Table 4: Comparison of scaling-based quantization methods.
As shown in Table 4, SmoothQuant [6] uses calibration to compute $S$. AWQ [11], as a weight-only quantization, approaches the problem from the perspective of protecting salient weights. Its scaling parameter is determined by the average magnitude of the activation, while the optimal exponent $\alpha^{*}$ is found by minimizing the loss through grid searching. Outlier Suppression+ [7] also uses grid searching optimization to find the optimal scaling after applying shifting. Unlike adopting the hand-crafting parameters in calibration, Outlier Suppression+ [7] uses grid searching optimization to find the optimal scaling after applying shifting. OmniQuant [2] optimizes the quantization parameters via gradient-based methods to minimize the block-wise quantization error. FlatQuant [12] and OSTQuant [4] both use learnable per-channel scaling, but at different stages of pre-quantization transformation, leading to different practical meanings.
Rotation
The rotation operation is a technique that adjusts data distribution by multiplying the data matrix by an orthogonal matrix. This operation largely changes the activation distributions, reduces outliers in the data, and thereby enhances quantization accuracy. The rotation process in a linear layer can be formulated as:
@@
Methods Rotation Strategy Rotation parameters $O$
QuIP [9] Rotation only Random orthogonal matrices
QuIP# [10] Rotation only Random Hadamard matrices
QuaRot [13] Rotation only Random Hadamard matrices
SpinQuant [3] Rotation only Gradient-based optimization
FlatQuant [12] Post-scaling Rotation Gradient-based optimization
OSTQuant [4] Pre-scaling Rotation Gradient-based optimization
Table 5: Comparison of rotation-based quantization methods.
@@
As shown in Table 5, QuIP [9] first introduced rotation with random orthogonal matrices to reduce incoherence in weight-only quantization. QuIP# [10] replaces this with random Hadamard matrices for greater efficiency. QuaRot [13] extends rotation to both weights and activations quantization using random Hadamard matrices, effectively eliminating outliers. SpinQuant [3] further improves stability by optimizing the rotation matrix on the Stiefel manifold to minimize quantization-induced performance variance, while still defaulting to random Hadamard for a few online rotations.
@@
Rotation is a highly effective pre-quantization transformation technique for reducing outliers in low-bit quantization, and it has also been widely adopted in many other quantization methods together with other pre-quantization transformation techniques to further improve performance (e.g., ResQ [17], QServe [15], FlatQuant [12], and OSTQuant [4], etc).
2.1.2 Quantization error mitigation
@@
The quantization error mitigation step employs techniques that mitigate the error introduced by quantization by compensating for it. This error between the original model and the quantized model can be computed from calibration datasets, and compensated to the original weights in linear layers or through low-rank branches. Fig. 2 demonstrates two commonly used quantization error mitigation methods.
Figure 2: Quantization error mitigation: compensate the error produced by quantization. Here, $dq(W_{q})$ denotes dequantizing $W_{q}$ back to the original precision value space, for computational purposes.
Self-compensation
We define self-compensation as directly compensating for errors on the original weights. Existing methods that employ self-compensation are generally based on analyzing the loss error introduced by quantization and make use of its Hessian-based quadratic approximation, which can be formulated as:
$\Delta\mathcal{L}=\mathcal{L}(dq(W_{q}))-\mathcal{L}(W)\approx\frac{1}{2}(dq(W_{q})-W)^{\top}H(dq(W_{q})-W)$ (4)
where $\mathcal{L}(W)$ is the loss function of the model evaluated at weights $W$. $W_{q}$ and $dq(W_{q})$ denote the quantized weights and their dequantized version, respectively. Self-compensation methods need to find an optimal weight perturbation $\delta$ to add to $W$, such that the quantization loss error in Eq. (4) is minimized. Table 6 summarizes the computation strategies of $\delta$ adopted by various self-compensation quantization methods.
Methods Weight perturbation $\delta$ computation
OBQ [18] Layer-wise MSE
GPTQ [8] Layer-wise MSE
QuIP [9] LDL decomposition
GPTAQ [19] Consider input deviation
GQuant [20] Add gradient information
Table 6: Comparison of self-compensation quantization methods.
Some methods simplify the quantization problem to minimize the squared error between the quantized and original output of each layer [18, 8]. Based on the squared error loss function, OBQ [18] and GPTQ [8] compute the Hessian matrix used in Eq. (4) and further calculate the optimal weight perturbation $\delta$. Unlike OBQ, which uses greedy orders to quantize weights, GPTQ quantizes all rows of weights in a fixed order to achieve a faster quantization process.
Some other methods take into account additional factors beyond the linear layer-wise formulation. While also aiming to minimize the quadratic proxy of the weight error in Eq. (4), QuIP [9] performs an LDL decomposition of the Hessian matrix to construct a compensation term that corrects for quantization-induced errors. GPTAQ [19], built upon GPTQ, addresses the limitation that GPTQ’s layer-wise optimization objective ignores the deviation between a layer’s actual input and that under the full-precision model. GQuant [20] further observes that the layer-wise MSE treats all hidden features equally, and therefore incorporates per-feature gradients into the objective function to improve quantization accuracy.
Low-rank decomposition compensation
@@
Beyond compensating errors directly on the original weights, an alternative approach is to introduce an auxiliary branch dedicated to error correction.
Some prior works leverage a low-rank term to reconstruct quantization error (e.g., ZeroQuant-v2 [14], CALDERA [21], LQER [16], QERA [22], etc), where the low-rank branch can either be quantized to higher precision, such as INT8, or remain in FP16. After incorporating a low-rank branch, the output $Y$ is computed as:
$Y=Xdq(W_{q})+XAB$ (5)
@@
where $A\in\mathbb{R}^{C_{in}\times r}$, $B\in\mathbb{R}^{r\times C_{out}}$ are low-rank matrices, and $r\ll\min(d,k)$ is the bottleneck rank. The product $AB\in\mathbb{R}^{d\times k}$ serves as a low-rank approximation of the quantization error. Table 7 illustrates the computation process of the low-rank terms in some quantization methods.
Methods Low rank terms $A,B$ computation
ZeroQuant-v2 [14] $\text{SVD}(dq(W_{q})-W)$
LQER [16] $\text{SVD}(dq(W_{q})-W)$
L2QER [16] $\text{SVD}(S(dq(W_{q})-W)),A:=S^{-1}A$
CALDERA [21] LPLR factorization [21]
QERA [22] $\text{SVD}(R^{\frac{1}{2}}_{X\!X}(dq(W_{q})-W)),A:=(R^{\frac{1}{2}}_{X\!X})^{-1}A$
Table 7: Comparison of low-rank decomposition compensation quantization methods. $A,B$ represents the low rank terms. $R_{X\!X}$ denotes the autocorrelation matrix of the input $X$.
@@
ZeroQuant-v2 [14] and LQER [16] apply SVD composition to reconstruct the quantization error matrix $E=W-W_{q}$. Based on LQER, L2QER, which is proposed in LQER, scales the $E$ before applying SVD heuristically, and inversely applies scales in $A_{k}$. This operation aims to more precisely approximate the quantization error corresponding to the salient weights.
In addition to the aforementioned approaches, CALDERA [21] and QERA [22] derive the optimal low-rank terms for quantization error mitigation by framing it as an optimization problem that minimizes the squared error between each layer’s output in Eq. (5) and its original output.
@@
In contrast to approaches that introduce low-rank branches externally, QwT [23] integrates lightweight linear layers within each decoder layer block of LLMs, directly compensating quantization errors, offering new insights for quantization error mitigation.
2.2 Settings of quantization
2.2.1 Symmetric vs. asymmetric